how much should a food manufacturer sell pound of canned meat if his cost
is 30 pesos per can and he wants to make a 20% profit on the selling price
for after giving away an 8% discount?

Answer by Edwin McCravy(20054)

how much should a food manufacturer sell pound of canned meat if
his cost is 30 pesos per can and he wants to make a 20% profit on
the selling price for after giving away an 8% discount?
Let the selling price be s.

He wants to reduce the selling price by 8% leaving only 92% of s, 
or 0.92s,

which is what he will receive.

He wants this 0.92s to be his cost of 30 pesos plus 20% of his cost 
of 30 pesos.

That's 30 + (0.20)(30) = 30+6 = 36 pesos.

So he wants 0.92s to equal 36 pesos.

0.92s = 36
 
s = 36/0.92 = 39.13043478... pesos

Round that to the nearest centavo, 39.13 pesos.  <-- answer

We check.

He begins with 39.13 pesos.

He reduce it by 8%, and 0.08(39.14) = 3.1384 pesos, rounded 
to 3.14 pesos

So he sells it for 39.3-3.14 = 36.16 pesos

Since he paid 30 pesos, his profit is 6.16 person

Now we find out what percent of 30 pesos his profit of 6.16 pesos is.

So we calculate 6.16/30 = .20533333... which rounds to 21%.

The difference is due to the fact that we must round off.

Edwin